# Inference of Human-derived Specifications of Object Placement via Demonstration

### v1

@@ 1 Introduction
As robots become a mainstay of industrial and manufacturing processes, pick-and-place tasks (e.g., packing, sorting, and kitting objects) have become central to many of their applications Sanneman et al. (2020). While significant prior research has explored control and manipulation for pick-and-place tasks, less has examined how robots can understand human preferences about object arrangement. Erbayrak et al, for example, designed an algorithm to pack objects into multiple bins while optimizing to keep as many objects of the same predefined “family” together as possible Erbayrak et al. (2021). Sun et al studied algorithms designed to instruct warehouse workers where to place items in a box, and observed when workers knowingly deviated from the algorithm’s plan Sun et al. (2022). The researchers then proposed modified algorithms to minimize human deviation from plans. While such approaches show promise in specific domains, they have limited expressiveness for required spatial relationships during object placement tasks. Therefore, in this work we present Positionally-Augmented Region Connection Calculus (PARCC), a spatial specification language able to capture humans’ requirements during object placement tasks. Additionally, we introduce an inference algorithm to infer PARCC specifications from demonstrations.
@@
Broadly speaking, capturing humans’ understanding of objects’ spatial relationships in a scene is not a new topic to research. Paul et al introduced a framework to “ground” human instructions in a world representation (e.g., understanding the instruction “pick up the middle block in the row of 5 blocks”) Paul et al. (2018), while others have developed methods to generate scenes or object arrangements from descriptions, either via predefined propositions (e.g., “A is left of B”) Wiebrock et al. (2000) or through natural language Vasardani et al. (2013); Liu et al. (2022).
While these systems could theoretically be used by workers to communicate specifications, there are a few underlying limitations. First, these methods only describe one scene at a time: for example, in Figure 1, prior methods may be able to express that the can $c_{2}$ is the furthest-north object in its scene, or generate an approximation of a scene via description of each object’s placement. However, when considering a specification describing both scenes, these methods cannot communicate concepts such as “all cans are east of oranges” or “all oranges must touch another orange to the north or south.” Capturing these relational rules can ensure, for example, that more fragile objects like apples or oranges are properly supported via contact or that there is consistent spatial placing of objects that may be expected by human workers. Additionally, methods capturing a human’s understanding of a scene rely upon the human directly providing a description of that scene Wiebrock et al. (2000); Vasardani et al. (2013); Liu et al. (2022). However, humans often under-specify (or misspecify) tasks, relative to what robots require, leading to undesired behavior when following a human’s directly-provided specification Gross et al. (2016).
@@ 2 Related Works
Within the field of task specification via formal logics, descriptions of spatial specifications often use Signal Temporal Logic (STL), since its operation over continuous signals makes it ideal for expressing spatial preferences Maler and Nickovic (2004). Application of STL to spatial problems either uses standard STL operators to express positions and regions as signals Linard and Tumova (2020) or modifies notation to include spatial-specific operators Nenzi et al. (2015); Ma et al. (2020). For example, Nenzi et al added two spatial modalities (“somewhere” and “surrounds”) to STL that operate over an undirected graph representing space (e.g., $\phi_{1}\mathcal{S}_{[d_{1},d_{2}]}\phi_{2}$ expresses that a region where $\phi_{1}$ is true is surrounded by the region where $\phi_{2}$ is true) Nenzi et al. (2015).
Other spatial specification languages use quad-tree representations Haghighi et al. (2015). Here, space is recursively partitioned into quadrants over which a specification reasons. For example, a quad-tree may represent a city and the specification describes power grid requirements across the city.
While STL and quad-tree representations are highly expressive, they do not lend themselves to cleanly encoding human-intuitive spatial relationships between objects, which tend to use qualitative descriptors and small, countable values. Conversely, region connection calculus (RCC) is a spatial-relational language introduced by Randel et al to formalize human-intuitive concepts of spatial relationships between regions Randell et al. (1992). The fundamental relation of RCC is $C(x,y)$ — read as ‘$x$ connects with $y$’, meaning that the topological closure of regions $x$ and $y$ share at least one point. The two axioms for C are as follows:
@@
While RCC itself is not based in logical specification, it has been adopted into specification languages. Ven et al introduced qualitative privacy description language (QPDL), using RCC within linear temporal logic (LTL) to describe technological privacy van de Ven and Dylla (2016). Similarly, spatio-temporal synthesis logic (STSL) combines $SU_{4}$ (a spatial language similar to RCC) with STL to characterize spatio-temporal dynamic behaviors in applications such as adaptive cruise control Li et al. (2020). However, neither QPDL nor STSL can express specification over “classes” of regions (i.e., “oranges are east of cans,” as shown in Figure 1), nor has an algorithm been proposed to infer such specifications from demonstration.
With respect to specification inference via demonstrations, we take inspiration from Vazquez et al, who proposed an inference framework over LTL via a maximum entropy approach Vazquez-Chanlatte et al. (2018). This framework derives a likelihood model over specifications based on how frequently demonstrations satisfy a specification versus the probability that the specification would be satisfied by random actions.
@@ 3 PARCC Formulation
We constrain PARCC to exist over axis-aligned rectangular objects on a flat plane; this choice is motivated by the ubiquitous use of rectangular bounding boxes to represent the size and location of objects in a scene Ali and Zhang (2024); Zendehdel et al. (2023); He et al. (2021); Jia et al. (2021). In this paper, we define an object as a tuple, $o=(o_{l},o_{w},o_{x},o_{y},o_{c})$, where $o_{l}$ is the object’s length, $o_{w}$ is its height, $o_{x}$ and $o_{y}$ designate its position in space, and $o_{c}$ is the object’s “class” — a label provided to objects subject to the same specifications (e.g., apples, oranges, and cans in Figure 1). In application, this could designate fragile vs. non-fragile objects, shipping destinations for packages, etc.
@@ 3.1.1 PARCC Formulas
Prior work has used RCC relations as propositions in logic languages van de Ven and Dylla (2016); we extend this to Boolean logic over PARCC class relations.
@@ 4 Specification Inference
1 Function FindDisjunctions($D_{1}...D_{d}$, $\mathcal{L}$, NN):
@@ 4.1 Finding Satisfying Disjunctive Formulas
Our inference procedure begins by finding a set of disjunctive PARCC formulas, $\bar{\mathcal{C}}$, which satisfy all demonstrations, $\mathcal{D}$ (Algorithm 1). We use an exhaustive search-based method to determine $\bar{\mathcal{C}}$. While one can brute-force a search over all possible disjunctive formulas, this would require checking $\mathcal{O}(|\mathcal{L}|^{2N}4^{N})$ formulas. Therefore, we take inspiration from Shah et al., and allow the use of production rule “templates” limiting the search space based on domain knowledge (see Section 5.2 for an example) Shah et al. (2018).
@@ 4.2 Determining Intended Disjunctive Formulas
1 Function Inference($\mathcal{D}$, $\bar{\mathcal{C}}$, $p_{c}$, $k_{r}$):
@@ 5.1 Experimental Setup
First, the subject provided eight demonstrations, $\mathcal{D}_{D}$, attempting to follow all spatial patterns they observed in $\mathcal{D}_{I}$. (Note we had not yet introduced the PARCC language to participants, and they were free to consider “patterns” in whatever representation was most natural.) The participant then provided a natural-language explanation of spatial patterns in $\mathcal{D}_{D}$ (again, the form of these “patterns” were determined entirely by the participant). Finally, we introduced participants to the PARCC language, and had them provide a PARCC specification $\Phi_{S}$ that best described their demonstrations, $\mathcal{D}_{D}$. Our algorithm then generated two demonstrations: $C_{D}$, which optimized satisfaction of the specification $\Phi_{D}$ inferred from $\mathcal{D}_{D}$; and $C_{S}$, which optimized the specification $\Phi_{S}$ provided directly by the user. These computer-generated demonstrations allowed us to probe subjects’ response to system behavior when following specifications inferred from demonstrations ($\Phi_{D}$) versus those explicitly specified ($\Phi_{S}$). To generate $C_{D}$ and $C_{S}$, we used a combination of mixed-integer programming and Monte Carlo tree search, both of which have been employed in box-packing domains Erbayrak et al. (2021); Edelkamp et al. (2014). Finally, we showed the participant both computer-generated demonstrations simultaneously, $C_{D}$ on the left and $C_{S}$ on the right, and asked the following Likert-style questions:
@@ 5.2 Choice of Search Template
Similar to Dwyer’s identification of LTL templates applicable to many real-world tasks, this template is intended to limit the search space to the most natural “constraints” for human operators across domains Dwyer et al. (1999). While a full investigation of template choice can constitute its own work, we compare our choice of template to two others in Technical Appendix D. One template is less restrictive than the one we use (i.e. allowing a larger set of possible disjunctive formulas) and one is more restrictive (i.e. allowing a smaller set of disjunctive formulas). The specifications inferred from human demonstrations ($\Phi_{D}$) are identical between our choice of template and the less restrictive template, showing that our template imposes minimal inductive bias on the specifications inferred in this experiment, and provides evidence that our template represents rules relevant to human demonstrators more generally. Additionally, we compare our template to a more restrictive template to show how operators can intentionally eliminate rules irrelevant to a particular domain.
@@ 5.5 Discussion
The statistical support for $H1$ and $H2$ confirms that obtaining human PARCC specification via demonstrations provides an advantage over direct specification. Additionally, support for $H4$ suggests the perceived difference between $C_{D}$ and $C_{S}$ results from human misspecification, even when specification is performed via a formal language with inherent semantics translatable into natural language. This result is not surprising, given prior research demonstrates that humans often have difficulty interpreting and providing specifications in formal specification languages, with or without translation to natural language Loomes and Vinter (1997); Vinter et al. (1996); Vinter (1998); Greenman et al. (2023).
@@ 6 Limitations and Future Work
While PARCC and the inference method in this paper captures spatial specifications relevant to human demonstrators, there are several open areas for future work. First, PARCC is limited to rectangular objects in two dimensions. Extending to three dimensions and more varied geometries can improve applicability in situations where objects cannot be treated as bounding boxes in an image. However, considering a third dimension requires modeling object stacking and stability, which is outside this paper’s scope. Second, the inference procedure relies on search over all possible disjunctive formulas allowed by a template. While this works for a few object classes, it may be intractable for a larger number of classes due to combinatorial explosion. In these cases, a sample based approach similar to Shah et al’s use of Markov Chain Monte Carlo may be useful Shah et al. (2018). Finally, inference over PARCC only considers rules that exist over all demonstrations. In many situations, learning preferences in addition to strict rules may provide a route to capture more nuanced aspects of human demonstrations.

### v2

@@ 1 Introduction
As robots become a mainstay of industrial and manufacturing processes, pick-and-place tasks (e.g., packing, sorting, and kitting objects) have become central to many of their applications state-of-robots. While significant prior research has explored control and manipulation for pick-and-place tasks, less has examined how robots can understand human preferences about object arrangement. Erbayrak et al, for example, designed an algorithm to pack objects into multiple bins while optimizing to keep as many objects of the same predefined “family” together as possible erbayrak2021multi. Sun et al studied algorithms designed to instruct warehouse workers where to place items in a box, and observed when workers knowingly deviated from the algorithm’s plan sun2022predicting. The researchers then proposed modified algorithms to minimize human deviation from plans. While such approaches show promise in specific domains, they have limited expressiveness for required spatial relationships during object placement tasks. Therefore, in this work we present Positionally-Augmented Region Connection Calculus (PARCC), a spatial specification language able to capture humans’ requirements during object placement tasks. Additionally, we introduce an inference algorithm to infer PARCC specifications from demonstrations.
@@
Broadly speaking, capturing humans’ understanding of objects’ spatial relationships in a scene is not a new topic to research. Paul et al introduced a framework to “ground” human instructions in a world representation (e.g., understanding the instruction “pick up the middle block in the row of 5 blocks”) paul2018efficient, while others have developed methods to generate scenes or object arrangements from descriptions, either via predefined propositions (e.g., “A is left of B”) wiebrock2000inference or through natural language vasardani2013descriptions; liu2022structformer.
While these systems could theoretically be used by workers to communicate specifications, there are a few underlying limitations. First, these methods only describe one scene at a time: for example, in Figure 1, prior methods may be able to express that the can $c_{2}$ is the furthest-north object in its scene, or generate an approximation of a scene via description of each object’s placement. However, when considering a specification describing both scenes, these methods cannot communicate concepts such as “all cans are east of oranges” or “all oranges must touch another orange to the north or south.” Capturing these relational rules can ensure, for example, that more fragile objects like apples or oranges are properly supported via contact or that there is consistent spatial placing of objects that may be expected by human workers. Additionally, methods capturing a human’s understanding of a scene rely upon the human directly providing a description of that scene wiebrock2000inference; vasardani2013descriptions; liu2022structformer. However, humans often under-specify (or misspecify) tasks, relative to what robots require, leading to undesired behavior when following a human’s directly-provided specification gross2016multi.
@@ 2 Related Works
Within the field of task specification via formal logics, descriptions of spatial specifications often use Signal Temporal Logic (STL), since its operation over continuous signals makes it ideal for expressing spatial preferences maler2004monitoring. Application of STL to spatial problems either uses standard STL operators to express positions and regions as signals linard2020active or modifies notation to include spatial-specific operators nenzi2015qualitative; ma2020sastl. For example, Nenzi et al added two spatial modalities (“somewhere” and “surrounds”) to STL that operate over an undirected graph representing space (e.g., $\phi_{1}\mathcal{S}_{[d_{1},d_{2}]}\phi_{2}$ expresses that a region where $\phi_{1}$ is true is surrounded by the region where $\phi_{2}$ is true) nenzi2015qualitative.
Other spatial specification languages use quad-tree representations haghighi2015spatel. Here, space is recursively partitioned into quadrants over which a specification reasons. For example, a quad-tree may represent a city and the specification describes power grid requirements across the city.
While STL and quad-tree representations are highly expressive, they do not lend themselves to cleanly encoding human-intuitive spatial relationships between objects, which tend to use qualitative descriptors and small, countable values. Conversely, region connection calculus (RCC) is a spatial-relational language introduced by Randel et al to formalize human-intuitive concepts of spatial relationships between regions randell1992spatial. The fundamental relation of RCC is $C(x,y)$ — read as ‘$x$ connects with $y$’, meaning that the topological closure of regions $x$ and $y$ share at least one point. The two axioms for C are as follows:
@@
While RCC itself is not based in logical specification, it has been adopted into specification languages. Ven et al introduced qualitative privacy description language (QPDL), using RCC within linear temporal logic (LTL) to describe technological privacy van2016qualitative. Similarly, spatio-temporal synthesis logic (STSL) combines $SU_{4}$ (a spatial language similar to RCC) with STL to characterize spatio-temporal dynamic behaviors in applications such as adaptive cruise control li2020stsl. However, neither QPDL nor STSL can express specification over “classes” of regions (i.e., “oranges are east of cans,” as shown in Figure 1), nor has an algorithm been proposed to infer such specifications from demonstration.
With respect to specification inference via demonstrations, we take inspiration from Vazquez et al, who proposed an inference framework over LTL via a maximum entropy approach vazquez2018learning. This framework derives a likelihood model over specifications based on how frequently demonstrations satisfy a specification versus the probability that the specification would be satisfied by random actions.
@@ 3 PARCC Formulation
We constrain PARCC to exist over axis-aligned rectangular objects on a flat plane; this choice is motivated by the ubiquitous use of rectangular bounding boxes to represent the size and location of objects in a scene ali2024yolo; zendehdel2023real; he2021know; jia2021self. In this paper, we define an object as a tuple, $o=(o_{l},o_{w},o_{x},o_{y},o_{c})$, where $o_{l}$ is the object’s length, $o_{w}$ is its height, $o_{x}$ and $o_{y}$ designate its position in space, and $o_{c}$ is the object’s “class” — a label provided to objects subject to the same specifications (e.g., apples, oranges, and cans in Figure 1). In application, this could designate fragile vs. non-fragile objects, shipping destinations for packages, etc.
@@ 3.1.1 PARCC Formulas
Prior work has used RCC relations as propositions in logic languages van2016qualitative; we extend this to Boolean logic over PARCC class relations.
@@ 4 Specification Inference
1Function FindDisjunctions($D_{1}...D_{d}$, $\mathcal{L}$, NN):
@@ 4.1 Finding Satisfying Disjunctive Formulas
Our inference procedure begins by finding a set of disjunctive PARCC formulas, $\bar{\mathcal{C}}$, which satisfy all demonstrations, $\mathcal{D}$ (Algorithm 1). We use an exhaustive search-based method to determine $\bar{\mathcal{C}}$. While one can brute-force a search over all possible disjunctive formulas, this would require checking $\mathcal{O}(|\mathcal{L}|^{2N}4^{N})$ formulas. Therefore, we take inspiration from Shah et al., and allow the use of production rule “templates” limiting the search space based on domain knowledge (see Section 5.2 for an example) shah2018bayesian.
@@ 4.2 Determining Intended Disjunctive Formulas
1Function Inference($\mathcal{D}$, $\bar{\mathcal{C}}$, $p_{c}$, $k_{r}$):
@@ 5.1 Experimental Setup
First, the subject provided eight demonstrations, $\mathcal{D}_{D}$, attempting to follow all spatial patterns they observed in $\mathcal{D}_{I}$. (Note we had not yet introduced the PARCC language to participants, and they were free to consider “patterns” in whatever representation was most natural.) The participant then provided a natural-language explanation of spatial patterns in $\mathcal{D}_{D}$ (again, the form of these “patterns” were determined entirely by the participant). Finally, we introduced participants to the PARCC language, and had them provide a PARCC specification $\Phi_{S}$ that best described their demonstrations, $\mathcal{D}_{D}$. Our algorithm then generated two demonstrations: $C_{D}$, which optimized satisfaction of the specification $\Phi_{D}$ inferred from $\mathcal{D}_{D}$; and $C_{S}$, which optimized the specification $\Phi_{S}$ provided directly by the user. These computer-generated demonstrations allowed us to probe subjects’ response to system behavior when following specifications inferred from demonstrations ($\Phi_{D}$) versus those explicitly specified ($\Phi_{S}$). To generate $C_{D}$ and $C_{S}$, we used a combination of mixed-integer programming and Monte Carlo tree search, both of which have been employed in box-packing domains erbayrak2021multi; edelkamp2014monte. Finally, we showed the participant both computer-generated demonstrations simultaneously, $C_{D}$ on the left and $C_{S}$ on the right, and asked the following Likert-style questions:
@@ 5.2 Choice of Search Template
Similar to Dwyer’s identification of LTL templates applicable to many real-world tasks, this template is intended to limit the search space to the most natural “constraints” for human operators across domains dwyer1999patterns. While a full investigation of template choice can constitute its own work, we compare our choice of template to two others in Technical Appendix D. One template is less restrictive than the one we use (i.e. allowing a larger set of possible disjunctive formulas) and one is more restrictive (i.e. allowing a smaller set of disjunctive formulas). The specifications inferred from human demonstrations ($\Phi_{D}$) are identical between our choice of template and the less restrictive template, showing that our template imposes minimal inductive bias on the specifications inferred in this experiment, and provides evidence that our template represents rules relevant to human demonstrators more generally. Additionally, we compare our template to a more restrictive template to show how operators can intentionally eliminate rules irrelevant to a particular domain.
@@ 5.5 Discussion
The statistical support for $H1$ and $H2$ confirms that obtaining human PARCC specification via demonstrations provides an advantage over direct specification. Additionally, support for $H4$ suggests the perceived difference between $C_{D}$ and $C_{S}$ results from human misspecification, even when specification is performed via a formal language with inherent semantics translatable into natural language. This result is not surprising, given prior research demonstrates that humans often have difficulty interpreting and providing specifications in formal specification languages, with or without translation to natural language loomes1997formal; vinter1996seven; vinter1998evaluating; greenman2022little.
@@ 6 Limitations and Future Work
While PARCC and the inference method in this paper captures spatial specifications relevant to human demonstrators, there are several open areas for future work. First, PARCC is limited to rectangular objects in two dimensions. Extending to three dimensions and more varied geometries can improve applicability in situations where objects cannot be treated as bounding boxes in an image. However, considering a third dimension requires modeling object stacking and stability, which is outside this paper’s scope. Second, the inference procedure relies on search over all possible disjunctive formulas allowed by a template. While this works for a few object classes, it may be intractable for a larger number of classes due to combinatorial explosion. In these cases, a sample based approach similar to Shah et al’s use of Markov Chain Monte Carlo may be useful shah2018bayesian. Finally, inference over PARCC only considers rules that exist over all demonstrations. In many situations, learning preferences in addition to strict rules may provide a route to capture more nuanced aspects of human demonstrations.